Changing the Cost Formula for a Month to the Cost Formula for a Year

During the past year, the high and low use of three different resources for Fly High Airlines occurred in July and April. The resources are airplane depreciation, fuel, and airplane maintenance. The number of airplane flight hours is the driver. The total costs of the three resources and the related number of airplane flight hours are as follows:


Resource
Airplane Flight
Hours

Total Cost
Airplane depreciation:
High 44,000 $ 18,900,000
Low 28,000 $ 18,900,000
Fuel:
High 44,000 445,896,000
Low 28,000 283,752,000
Airplane maintenance:
High 44,000 16,096,000
Low 28,000 11,912,000

Required:

If required, round calculations and answers to nearest dollar.

1. Develop an annual cost formula for airplane depreciation.

Total annual cost of airplane depreciation =  x ( $ )

Develop an annual cost formula for fuel.

Total annual cost of fuel = $ × Annual number of airplane flight hours

Develop an annual cost formula for airplane maintenance.

Total annual cost of airplane maintenance = ( x $ ) + ( $ × Annual number of airplane flight hours  )

2. Using the three annual cost formulas that you developed, predict the cost of each resource in a year with 487,000 airplane flight hours.

Total annual cost of airplane depreciation $
Total annual cost of fuel $
Total annual cost of airplane maintenance
Answer #1
Answer
Answer 2 Total annual cost of airplane depreciation = 12*$ 18,800,000
Fuel variable cost = 445896000/44000= $10,134
Total annual cost of fuel = 10134*Annual no. of airplane flight hours
Airplane maintenance variable cost = (16544000-11784000)/(44000-28000)= $                298.5
Airplance maintenance fixed cost =16544000-(44000*297.5)= $         34,54,000
Annual cost formula =(12*3454000)+(297.5*Annual no. of airplane flight hours)
Answer 2 Total annual cost of airplane depreciation=12*18800000= $    22,56,00,000
Total annual cost of fuel=10134*487000= $ 4,93,52,58,000
Total annual cost of airplane maintenance=(12*3454000)+(297.5*487000)= $    18,63,30,500
Please hit LIKE button if this helped.